# Digital-Analog Quantum Computing with Qudits

### v1

@@ I Introduction
Quantum computation originated from the idea of simulating quantum systems with controllable quantum simulators [Benioff1980, Feynman1982]. Firstly, it was proposed a paradigm called analog quantum computing (AQC) [Das2008], in which a controllable quantum system, called a quantum simulator, emulates the dynamics of the quantum system of interest. Afterwards, a computational paradigm which proceeds through sequences of discrete quantum gates acting on one or more qubits was proposed, named digital quantum computing (DQC) [Deutsch1995]. While AQC is generally more robust to noise [Poggi2020, Altman2021, adrian2024, Fauseweh2024], it lacks the versatility of DQC, whose gate-based formulation enables universal quantum computation. In contrast, the susceptibility of DQC to noise can be suppressed by employing quantum error correction codes [aharonov1996, Kitaev1997, Bultrini2022]. These codes, however, impose substantial resource overheads, as a single logical qubit requires multiple high-quality physical qubits. Although recent progress in hardware development is moving towards larger implementations with better quality qubits [Wintersperger2023, Bluvstein2024, Bravyi2024, strohm2024ionbasedquantumcomputinghardware, AbuGhanem2025], current platforms still operate within the noisy intermediate-scale quantum (NISQ) regime [Preskill2018NISQ].
@@
To improve the performance of quantum processors in the NISQ era, a hybrid framework combining the robustness of AQC and the flexibility of DQC, known as digital-analog quantum computing (DAQC) [Parra2020], was proposed. DAQC integrates both paradigms by employing the system’s natural two-body Hamiltonian as an entangling resource, while digital single-qubit rotations are used to conjugate the Hamiltonian to yield a universal set of operations [PhysRevA.65.040301, Bassler2023synthesisof, ArbitraryDAQC]. This allows for the implementation not only of quantum simulations of arbitrary two-body Hamiltonians, but also quantum algorithms [Ana2020QFT, Ana2022QAOA, Ana2023HHL]. Within the DAQC paradigm, two implementation schemes are typically considered: stepwise-DAQC (sDAQC), where the Hamiltonian is switched on and off between single-qubit rotations, and banged-DAQC (bDAQC), in which the Hamiltonian remains active while single-qubit gates are applied on top of the continuous dynamics. Although the latter introduces an intrinsic error, it is experimentally more feasible, since the interaction Hamiltonian is not switched on and off, and exhibits superior scalability with respect to the dominant error sources in current quantum processors [Garcia2024].
@@
Universal quantum computation can also be achieved using qudits through appropriate two-qudit gates or an entangling Hamiltonian and local unitaries [Nielsen2002, Bremner2005, Brennen2005, Brennen2006]. Encoding quantum information in $d$-level quantum systems, known as qudits, enhances the computational capabilities of near-term devices by reducing the number of entangling operations, which might lead to improved experimental implementations [Wang2020, Nikolaeva2024, Kiktengo2025, Gedik2015]. Qudits offer distinct advantages in quantum simulation, where they provide a natural platform for studying high-dimensional spin phenomena [Senko2015, Wang2023], enabling, for instance, qubit–qudit hybrid algorithms [Bocharov2017, Meth2025]. Qudit-based processors have already been experimentally demonstrated on superconducting platforms [Goss2022, Luo2023, Nguyen2024, Yu2025] and trapped-ion architectures [Chi2022, Ringbauer2022], among others.
@@
In Ref. [Nielsen2002], it was proven that universal quantum simulation can be achieved employing a fixed two-body entangling $N$-qudit Hamiltonian and local unitaries, but they provided no constructive method. Afterwards, in Ref. [Bremner2005], they provided a method to achieve it, but with a worse circuit complexity than DAQC.
@@
In this article, we aim to generalize the DAQC protocols [Parra2020, ArbitraryDAQC] to qudit systems. More specifically, we extend the DAQC framework to simulate a target two-body qudit Hamiltonian using another arbitrary two-body qudit Hamiltonian as a resource, together with local qudit unitaries. Our constructive method allows us to design DAQC circuits which employ polynomially fewer applications of single-qudit gates than previous protocols relying on isolating every coupling [Bremner2005].
@@ II Revisiting the DAQC protocol for qubit Hamiltonians
The protocol introduced in Ref. [ArbitraryDAQC] systematically constructs the matrix $M$ of Eq. (4) from the complete family of conjugations of $H_{\text{S}}$ under the set of local single-qubit operators. This family comprises $2^{n-1}$ operators in the $ZZ$ case and $4^{n}$ operators for arbitrary Hamiltonians. The matrix $M$ can be decomposed up to $n+1$ blocks, ranging from $M_{[0]}$, which consists of a single column of ones corresponding to conjugation by the identity, to $M_{[n]}$, where all qubits are simultaneously conjugated by different combinations of Pauli operators. It has been shown in Ref. [ArbitraryDAQC] that this construction guarantees the existence of a positive-time solution.
@@ III Generalizing the protocol to qudit Hamiltonians
We now extend the DAQC protocol of Ref. [ArbitraryDAQC] to arbitrary local dimension $d$, i.e., to qudit circuits. The generalization relies on the complete Weyl–Heisenberg operator basis, a generalized Clifford algebra defined as [Bertlmann2008]
@@
where $w=e^{2\pi i/d}$ and $n,m\in[0,d-1]$. Here, $W_{10}=Z$ and $W_{01}=X$ denote the phase- and position-shift operators, generalizing the Pauli matrices through the clock and shift formalism. The Weyl operators satisfy the following relations [Popp2024]:
@@ IV.1 Numerical results
For the simulations, we employ the banged-DAQC approach [Parra2020], in which single-qudit rotations are applied while the Hamiltonian remains active, rather than switching it on and off, as this is experimentally more feasible. The initial density matrix corresponds to a GHZ state, which is maximally entangled.
Regarding the parameters, the thermalization time is set to $T_{1}=100T$, while the duration of the single-qutrit gate $X$ is $\Delta t=10^{-2}T$ with a fidelity of $\sim 99.4\%$ [Ringbauer2022]. When solving the linear system, solutions with $t_{i}<4\Delta t$ are discarded, since these blocks would be shorter than the longest single-qutrit conjugation times.
In the digital case, two-qutrit controlled-phase gates are implemented using the unitary evolution from Ref. [Goss2022], with a fixed fidelity of $95\%$, while single-qutrit gates are applied as detailed in Appendix C.
@@ V Conclusions
The theory developed here expands the range of systems that can be efficiently simulated with digital-analog protocols, for instance those containing quadrupolar terms. Additionally, qudit DAQC would eliminate the need for multi-qubit mappings in bosonic simulations [PhysRevLett.129.160501, PhysRevResearch.6.013202] while showing higher resilience against noises. Consequently, this paves the way for simulating gauge theories requiring fewer quantum resources, an essential advantage in the NISQ era.
The generalization to qudits provides a comprehensive experimental framework to implement DAQC circuits employing an entangling Hamiltonian and only the single-qudit operators $\{X_{d},Z_{d}\}$. This makes this approach implementable in state-of-the-art platforms, such as $d$-level system ion-traps [Ringbauer2022].
@@ Acknowledgements.
The authors thank David Aguirre and Yongcheng Ding for insightful and helpful discussions and acknowledge financial support from OpenSuperQ+100 (Grant No. 101113946) of the EU Flagship on Quantum Technologies, from Project Grant No. PID2024-156808NB-I00, also from Spanish Ramón y Cajal Grant No. RYC-2020-030503-I funded by MICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future”. This project has also received support from the Spanish Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain, and by the EU through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda. We acknowledge funding from Basque Government through Grant No. IT1470-22, and the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. AA acknowledges grant PID2021-125823NA-I00 funded by MICIU/AEI/10.13039/501100011033 and by the “FSE+”.
@@ A.1 $M_{[2]}$ uniqueness
We now prove that the determinant of $M_{[2]}$ is nonzero, ensuring the existence of a solution. To this end, we turn to M. H. Ingraham’s note on determinants [APX:Ingraham1937], which states that if a square matrix $A=(A_{ij})$ is composed of square submatrices $A_{ij}$ that commute pairwise, then the determinant of the matrix constructed by taking $\det(A_{ij})$ as its elements is equal to $\det(A)$. For illustration, consider the $2\times 2$ case
@@
Hence, we evaluate $\det M_{[2]}$ through its mutually commuting submatrices (see Eq. 26a). Using Eqs. 26b and 26c, the cross terms can be grouped into the set of matrices $\{M^{2},M^{0},M^{2}M^{1.1},M^{2}M^{1.2}\}$, as all products are powers of these matrices, as in Ref. [ArbitraryDAQC]. Consequently, the determinant takes the form of Eq. (28),
@@ Appendix B Proof of existence of a positive time solution
Following this reasoning, we can analogously construct a polytope as in Ref. [ArbitraryDAQC] with the matrix columns as the vertices. Then, we can find a hypersphere centered at the origin and which is completely inside the polytope, allowing us to prove the existence of a positive solution to the problem.
@@ Appendix C DQC circuit to simulate $H_{\text{P}}(\theta)$
where this Hamiltonian and the parameters $\alpha$ can be found in Ref. [Goss2022]. If we conjugate it, we obtain the isotropic bilinear–biquadratic (BLBQ) Ising-like Hamiltonian:
@@
For the simulation, we assume that the parameters can be approximated in such a way that both values are achieved, and we fix the fidelity of this two-qutrit gate at $95\%$, inspired by the fidelity of the $CZ$ gate. To minimize error propagation, the circuit is constructed in a brick-wall configuration [Tepaske2024], as shown in Fig. 4.
@@ C.2.2 $M^{2}$ eigenvalues
Since $A$ and $B$ commute, we will rely on M. H. Ingraham’s note on determinants [Ingraham1937]. It states that if a square matrix $C=(C_{ij})$ consists of pairwise commuting square subblocks, then the determinant of the matrix built from $\det(C_{ij})$ coincides with $\det(C)$. The determinant is then

### v2

@@ I Introduction
Quantum computation originated from the idea of simulating quantum systems with controllable quantum simulators [7, 20]. Firstly, it was proposed a paradigm called analog quantum computing (AQC) [16], in which a controllable quantum system, called a quantum simulator, emulates the dynamics of the quantum system of interest. Afterwards, a computational paradigm which proceeds through sequences of discrete quantum gates acting on one or more qubits was proposed, named digital quantum computing (DQC) [17]. While AQC is generally more robust to noise [38, 5, 45, 19], it lacks the versatility of DQC, whose gate-based formulation enables universal quantum computation. In contrast, the susceptibility of DQC to noise can be suppressed by employing quantum error correction codes [4, 29, 15]. These codes, however, impose substantial resource overheads, as a single logical qubit requires multiple high-quality physical qubits. Although recent progress in hardware development is moving towards larger implementations with better quality qubits [48, 9, 11, 43, 3], current platforms still operate within the noisy intermediate-scale quantum (NISQ) regime [41].
@@
To improve the performance of quantum processors in the NISQ era, a hybrid framework combining the robustness of AQC and the flexibility of DQC, known as digital-analog quantum computing (DAQC) [37], was proposed. DAQC integrates both paradigms by employing the system’s natural two-body Hamiltonian as an entangling resource, while digital single-qubit rotations are used to conjugate the Hamiltonian to yield a universal set of operations [18, 6, 21]. This allows for the implementation not only of quantum simulations of arbitrary two-body Hamiltonians, but also quantum algorithms [32, 26, 31]. Within the DAQC paradigm, two implementation schemes are typically considered: stepwise-DAQC (sDAQC), where the Hamiltonian is switched on and off between single-qubit rotations, and banged-DAQC (bDAQC), in which the Hamiltonian remains active while single-qubit gates are applied on top of the continuous dynamics. Although the latter introduces an intrinsic error, it is experimentally more feasible, since the interaction Hamiltonian is not switched on and off, and exhibits superior scalability with respect to the dominant error sources in current quantum processors [22].
@@
Universal quantum computation can also be achieved using qudits through appropriate two-qudit gates or an entangling Hamiltonian and local unitaries [35, 12, 14, 13]. Encoding quantum information in $d$-level quantum systems, known as qudits, enhances the computational capabilities of near-term devices by reducing the number of entangling operations, which might lead to improved experimental implementations [46, 36, 28, 23]. Qudits offer distinct advantages in quantum simulation, where they provide a natural platform for studying high-dimensional spin phenomena [42, 47], enabling, for instance, qubit–qudit hybrid algorithms [10, 33]. Qudit-based processors have already been experimentally demonstrated on superconducting platforms [25, 30, 34, 49] and trapped-ion architectures [1, 2], among others.
@@
In Ref. [35], it was proven that universal quantum simulation can be achieved employing a fixed two-body entangling $N$-qudit Hamiltonian and local unitaries, but they provided no constructive method. Afterwards, in Ref. [12], they provided a method to achieve it, but with a worse circuit complexity than DAQC.
@@
In this article, we aim to generalize the DAQC protocols [37, 21] to qudit systems. More specifically, we extend the DAQC framework to simulate a target two-body qudit Hamiltonian using another arbitrary two-body qudit Hamiltonian as a resource, together with local qudit unitaries. Our constructive method allows us to design DAQC circuits which employ polynomially fewer applications of single-qudit gates than previous protocols relying on isolating every coupling [12].
@@ II Revisiting the DAQC protocol for qubit Hamiltonians
The protocol introduced in Ref. [21] systematically constructs the matrix $M$ of Eq. (4) from the complete family of conjugations of $H_{\text{S}}$ under the set of local single-qubit operators. This family comprises $2^{n-1}$ operators in the $ZZ$ case and $4^{n}$ operators for arbitrary Hamiltonians. The matrix $M$ can be decomposed up to $n+1$ blocks, ranging from $M_{[0]}$, which consists of a single column of ones corresponding to conjugation by the identity, to $M_{[n]}$, where all qubits are simultaneously conjugated by different combinations of Pauli operators. It has been shown in Ref. [21] that this construction guarantees the existence of a positive-time solution.
@@ III Generalizing the protocol to qudit Hamiltonians
We now extend the DAQC protocol of Ref. [21] to arbitrary local dimension $d$, i.e., to qudit circuits. The generalization relies on the complete Weyl–Heisenberg operator basis, a generalized Clifford algebra defined as [8]
@@
where $w=e^{2\pi i/d}$ and $n,m\in[0,d-1]$. Here, $W_{10}=Z$ and $W_{01}=X$ denote the phase- and position-shift operators, generalizing the Pauli matrices through the clock and shift formalism. The Weyl operators satisfy the following relations [40]:
@@ IV.1 Numerical results
For the simulations, we employ the banged-DAQC approach [37], in which single-qudit rotations are applied while the Hamiltonian remains active, rather than switching it on and off, as this is experimentally more feasible. The initial density matrix corresponds to a GHZ state, which is maximally entangled.
Regarding the parameters, the thermalization time is set to $T_{1}=100T$, while the duration of the single-qutrit gate $X$ is $\Delta t=10^{-2}T$ with a fidelity of $\sim 99.4\%$ [2]. When solving the linear system, solutions with $t_{i}<4\Delta t$ are discarded, since these blocks would be shorter than the longest single-qutrit conjugation times.
In the digital case, two-qutrit controlled-phase gates are implemented using the unitary evolution from Ref. [25], with a fixed fidelity of $95\%$, while single-qutrit gates are applied as detailed in Appendix C.
@@ V Conclusions
The theory developed here expands the range of systems that can be efficiently simulated with digital-analog protocols, for instance those containing quadrupolar terms. Additionally, qudit DAQC would eliminate the need for multi-qubit mappings in bosonic simulations [24, 39] while showing higher resilience against noises. Consequently, this paves the way for simulating gauge theories requiring fewer quantum resources, an essential advantage in the NISQ era.
The generalization to qudits provides a comprehensive experimental framework to implement DAQC circuits employing an entangling Hamiltonian and only the single-qudit operators $\{X_{d},Z_{d}\}$. This makes this approach implementable in state-of-the-art platforms, such as $d$-level system ion-traps [2].
@@ Acknowledgements.
The authors thank David Aguirre and Yongcheng Ding for insightful and helpful discussions and acknowledge financial support from OpenSuperQ+100 (Grant No. 101113946) of the EU Flagship on Quantum Technologies, from Project Grant No. PID2024-156808NB-I00, also from Spanish Ramón y Cajal Grant No. RYC-2020-030503-I funded by MICIU/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” and “ERDF Invest in your Future”, and from the Spanish Ministry of Science and Innovation under the Recovery, Transformation and Resilience Plan (CUCO, MIG-20211005). This project has also received support from the Spanish Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain, and by the EU through the Recovery, Transformation and Resilience Plan – NextGenerationEU within the framework of the Digital Spain 2026 Agenda. We acknowledge funding from Basque Government through Grant No. IT1470-22, and the IKUR Strategy under the collaboration agreement between Ikerbasque Foundation and BCAM on behalf of the Department of Education of the Basque Government. AA acknowledges grant PID2021-125823NA-I00 funded by MICIU/AEI/10.13039/501100011033 and by the “FSE+”.
@@ A.1 $M_{[2]}$ uniqueness
We now prove that the determinant of $M_{[2]}$ is nonzero, ensuring the existence of a solution. To this end, we turn to M. H. Ingraham’s note on determinants [27], which states that if a square matrix $A=(A_{ij})$ is composed of square submatrices $A_{ij}$ that commute pairwise, then the determinant of the matrix constructed by taking $\det(A_{ij})$ as its elements is equal to $\det(A)$. For illustration, consider the $2\times 2$ case
@@
Hence, we evaluate $\det M_{[2]}$ through its mutually commuting submatrices (see Eq. 26a). Using Eqs. 26b and 26c, the cross terms can be grouped into the set of matrices $\{M^{2},M^{0},M^{2}M^{1.1},M^{2}M^{1.2}\}$, as all products are powers of these matrices, as in Ref. [21]. Consequently, the determinant takes the form of Eq. (28),
@@ Appendix B Proof of existence of a positive time solution
Following this reasoning, we can analogously construct a polytope as in Ref. [21] with the matrix columns as the vertices. Then, we can find a hypersphere centered at the origin and which is completely inside the polytope, allowing us to prove the existence of a positive solution to the problem.
@@ Appendix C DQC circuit to simulate $H_{\text{P}}(\theta)$
where this Hamiltonian and the parameters $\alpha$ can be found in Ref. [25]. If we conjugate it, we obtain the isotropic bilinear–biquadratic (BLBQ) Ising-like Hamiltonian:
@@
For the simulation, we assume that the parameters can be approximated in such a way that both values are achieved, and we fix the fidelity of this two-qutrit gate at $95\%$, inspired by the fidelity of the $CZ$ gate. To minimize error propagation, the circuit is constructed in a brick-wall configuration [44], as shown in Fig. 4.
@@ C.2.2 $M^{2}$ eigenvalues
Since $A$ and $B$ commute, we will rely on M. H. Ingraham’s note on determinants [27]. It states that if a square matrix $C=(C_{ij})$ consists of pairwise commuting square subblocks, then the determinant of the matrix built from $\det(C_{ij})$ coincides with $\det(C)$. The determinant is then